# Scalable Quantum Simulations of Scattering in Scalar Field Theory on 120 Qubits

### v1

@@ I Introduction
Quantum simulation holds the potential for predictive power in a wide class of fundamental physics problems that are out of reach for classical methods. In particular, quantum simulations of high-energy inelastic scattering of relativistic particles, such as those that take place in extreme cosmological events or particle colliders, promise to shed light onto the structure and properties of the Standard Model and what may lie beyond [31]. Standard techniques in quantum field theory predict the outputs of scattering events in regimes where perturbation theory is valid. Although finite-volume lattice methods [32, 33] are able to compute non-perturbative quantities for specific processes (for example, see Refs. [34, 35, 36, 37, 38]), studying arbitrary scattering events is challenging. Even in the perturbative regime, conventional analytical methods only give static quantities, or results that are valid at infinitely late times. While this approximation is valid for many quantities of interest, the dynamics of collision events are generally inaccessible. Numerical lattice techniques suffer from an exponential increase of resources due to the sign problem when attempting to simulate real-time dynamics. Furthermore, classical simulations of highly-inelastic scattering using methods based on truncations of entanglement, such as Matrix Product States (MPS), are expected to fail because of the large amount of entanglement that is believed to be generated in these events [39, 40, 41, 42, 43, 44]. As such, the dynamics of scattering of fundamental particles is a prime example of a physically-relevant problem that is challenging for classical computation, making it a promising target for quantum simulation.
@@
Much work has been done toward the goal of full-scale quantum simulations of scattering in theories and regimes where analytic results are scarce and classical computations are intractable. Proposals of wavepacket creation techniques [45, 39, 46, 47, 48, 49] and strategies to extract scattering observables from dynamics [50, 51] have complemented demonstrations of quantum simulations of scattering in various models [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63]. One of the simplest relativistic field theories in which scattering may be studied is the scalar field theory in one dimension. [^1] While it has been proven that the scalar field theory is trivial in five spacetime dimensions, and believed to be trivial in four, the existence of a non-trivial continuum limit has been rigorously established in the one- and two-dimensional cases [64]. The continuum limit can be recovered from lattice simulations by studying systems with successively finer lattices and carefully tuning the system parameters as functions of the lattice spacing. This is the theory of coupled harmonic oscillators at each point in space with a quartic potential that governs interactions between particles. Scalar and pseudoscalar particles play key roles in many areas, including nuclear physics [65, 66], condensed matter physics [67, 68], high energy physics [69, 70, 71], as well as cosmology [72] and physics beyond the Standard Model [73, 74, 75]. Scalar fields have also been used as a venue to study phase transitions and spontaneous symmetry breaking [76, 77, 78]. Moreover, scattering in scalar field theory has been shown to be BQP-complete [79], meaning all “quantum-efficient” problems may be solved by mapping them to quantum simulations of scattering with various initial conditions. As such, this “simple” theory is a good sandbox for demonstrations of scalable quantum simulation methods. Following the seminal work of Jordan, Lee, and Preskill (JLP) [80, 81, 79], there have been numerous efforts targeting aspects of quantum simulations of scattering in scalar field theory [82, 83, 84, 85, 86, 87, 88, 89, 90, 91].
@@
In this work, variational algorithms on the basis of brickwall [92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102] circuits222“Brickwall”, “brickwork”, and “hardware-efficient” are commonly used to refer to dense variational ansatze using native single- and two-qubit gates. “Brickwall” is used to describe this class of circuits in this work. and other ansatze are used to simulate the scattering of two particle wavepackets on IBM’s quantum computers. While the application of variational algorithms to quantum simulations has been proposed and executed on devices in the past [103, 104, 105, 106, 107], this work serves as the first example of a quantum simulation using physics-informed brickwall circuits at scale for state preparation and time evolution. Furthermore, it is the first quantum simulation of scattering of relativistic particles. These proof-of-concept simulations are a demonstration of the techniques that will be foundational for simulations of high-energy, high-inelasticity scattering events in more complicated theories, where rich physical phenomena are expected. The simulations proceed in several steps:
@@
By incorporating the symmetries and physical structure of the scalar field theory into the circuit design, the simulation is carried out in a fully-scalable way, unlocking the full capabilities of state-of-the-art devices. These techniques guide the design of variationally-optimized circuits to minimize the circuit depth in order to maximize the utility of devices available today. While previous quantum simulations have taken a hybrid approach, (e.g., state preparation was done variationally, but time evolution was implemented using an algorithm with rigorous performance guarantees), this work establishes the application of physics-informed variational algorithms to all parts of the simulation. The variational methods developed in this work are expected to have broad relevance to simulating relevant physical models.
This paper is organized as follows. An overview of lattice scalar field theory in one dimension and its mapping to qubits is given in Sec. II. The algorithm for implementing scalable variational circuits (SVCs) for a class of quantum simulation problems is outlined in Sec. III. This algorithm may be seen as a general framework capable of compressing physical simulation circuits at scale. Sec. IV describes how knowledge of the physics of the theory is used to determine SVCs both for state preparation and time evolution. The errors stemming from these approximate circuit compression methods are quantified. In Sec. V, SVCs are used to simulate scattering in scalar field theory using 120 qubits of IBM’s superconducting quantum computer ibm_fez. Both the free and the interacting cases are studied. A new error mitigation strategy on the basis of simulations of vacuum evolution is introduced. With this technique, the results are found to be in qualitative agreement with classical MPS simulations, and the effect of the interaction strength on the nature of the collisions is clearly seen. Improvements and comments on the extensibility of the developed methods are presented in Sec. VI.
@@ II Lattice scalar field theory
$H$ $=\sum_{j=0}^{L-1}\frac{1}{2}\Pi_{j}^{2}+\frac{1}{2}m^{2}\phi_{j}^{2}+\frac{1}{2}(\phi_{j}-\phi_{j+1})^{2}+\frac{\lambda}{4!}\phi_{j}^{4}$ (1)
@@
Here $m$ is the bare mass, $\lambda$ is the coupling controlling the strength of $\phi^{4}$ interactions, $\Pi_{j}$ is the conjugate momentum operator obeying the canonical commutation relation $[\phi_{i},\Pi_{j}]=i\delta_{ij}$, and the nearest-neighbor finite-difference operator $\nabla\phi=(\phi_{j}-\phi_{j+1})$ is used in the lattice representation of the kinetic term. The lattice spacing is set to $a=1$ in this work.
With PBCs, the eigenstates of $H$ are labeled by their spatial momenta $k_{i}=0,\pm\frac{2\pi}{L},\pm\frac{4\pi}{L},\dots,+\pi$. The vector of fields $\vec{\phi}=(\phi_{0},\phi_{1},\dots,\phi_{L-1})$ is related to the vector of spatial momentum eigenstates $\vec{\chi}$ by a spatial Fourier transform: $\vec{\chi}=V\vec{\phi}$, where $V=\frac{1}{\sqrt{L}}e^{i\vec{k}\cdot\vec{j}}$ is the Fourier transform matrix. In the free theory $(\lambda=0)$, the energy of a single-particle eigenstate with spatial momentum $k$ is
@@
$H=\sum_{k}E_{k}\left(a_{k}^{\dagger}a_{k}+\frac{1}{2}\right).$ (4)
@@
Time-evolved observables in the free theory can be computed analytically by expanding the state in the appropriate basis of multiparticle excitations. For example, the time evolution of a single-wavepacket state can be written as $|\psi_{\text{wp}}\rangle(t)=\sum_{k}e^{-itE_{k}}g(k,p)a_{k}^{\dagger}|\psi_{\text{vac}}\rangle$. By rewriting $\phi^{2}_{j}$ in terms of $a_{k}$ and $a_{k}^{\dagger}$, $\langle\phi^{2}_{j}\rangle(t)$ can be obtained using Wick contractions. In the interacting $(\lambda\neq 0)$ theory, $E_{k},a_{k}^{\dagger},$ and $|\psi_{\text{vac}}\rangle$ are all functions of $\lambda$. In cases where $\frac{\lambda}{4!}$ is small, observables can be obtained by treating the $\phi^{4}$ term as a perturbation.
@@
[…]
where $T$ is the time-ordering operator. With a proper choice of $N$ and $t$ depending on the parameters $m$ and $\lambda$ [80, 81, 108], it can be shown that unwanted evolution can be sufficiently suppressed.
@@ II.1 Qubit representation
At each lattice site $j$, the field takes continuous values $\phi_{j}\in(-\infty,\infty)$. Since the Hilbert space at each $j$ is infinite, the field must be digitized to represent it on a quantum computer. The JLP digitization is used to map this system onto qubits [80, 81]. Given a register of $n_{q}$ qubits, a maximum magnitude $\phi_{\text{max}}$ is chosen, [^3] There exists an optimal choice of $\phi_{\text{max}}$ for a given $n_{q}$ as a result of the Nyquist-Shannon Sampling Theorem [109, 110, 111]. The optimal values of $\phi_{\text{max}}$ are determined numerically for small $n_{q}$ in Ref. [82], and analytically in Refs. [112, 113]. and the field at each point in space is evenly discretized in $2^{n_{q}}$ steps of size $\delta_{\phi}$ in the range $[-\phi_{\text{max}},\phi_{\text{max}}]$. At each site $j$, the digitized field can take on the values
@@
$k_{\phi,j}=-\frac{\pi}{\delta_{\phi}}+\left(\ell-\frac{1}{2}\right)\frac{2\pi}{2^{n_{q}}\delta_{\phi}},\quad\ell\in[0,2^{n_{q}}-1].$ (14)
The allowed state of the field at each site is represented by qubits using a binary encoding of the $2^{n_{q}}$ allowed values from Eqs. 13 and 14:
@@
The representation of the lattice scalar field by qubits is shown in Fig. 1. This construction efficiently recovers the low-energy physics in a way that is a systematically extendable to include higher-energy contributions. In Ref. [82] it was shown that, aside from device noise and errors stemming from approximations to the time evolution operator, the error in local observables decays double-exponentially with the amount of qubits used in the site-wise digitization. [^5] That is, for a local observable, the percent error $\epsilon$ in the low-energy eigenvalues of the digitized observable compared to their undigitized counterparts scales as $\epsilon\sim 2^{-2^{n_{q}}}$. In fact, it was found in Ref. [82] that $n_{q}=4$ provides sufficient precision in local observables, so that other sources of error dominate. Furthermore, it was also found that the $\phi$ basis is advantageous over other representations, such as the basis of harmonic oscillator eigenstates, and strategies associated with improved actions that are often used in lattice QCD. This advantage stems from the simplicity and efficiency of the implementations of operators in the theory (Eqs. (16), (17)).
The JLP prescription gives an efficient implementation of the operators using quantum gates [80, 81]. Operators in the Hamiltonian of Eq. (1) can be broken into two non-commuting groups: those that contain $\phi$, and those that contain $\Pi$. The operators made up of powers and tensor products of $\phi$ are diagonal in the $\phi$ basis, and are straightforward to implement. Aside from factors of the identity, $\phi^{2}_{j}$ and $(\phi_{j}-\phi_{j+1})^{2}$ contain only $ZZ$ operators and can be implemented using $O(n_{q}^{2})$ two-qubit gates. Similarly, $\phi^{4}_{j}$ can be implemented with $O(n_{q}^{4})$ two-qubit gates, as it has additional $ZZZZ$ terms for $n_{q}\geq 4$. $\Pi^{2}_{j}$ is translated into gates by using a site-wise Quantum Fourier Transform (QFT) to switch into the conjugate momentum basis where $\Pi$ is diagonal. Using this, $\Pi^{2}_{j}$ can be implemented with $O(n_{q}^{2})$ two-qubit gates as well. [^6] Note that for a device with nearest-neighbor (NN) two-qubit gate connectivity, as is the case for most superconducting quantum computers, there is an extra two-qubit gate overhead to convert circuits from arbitrary connectivity to NN. To reduce the two-qubit gate depth, efficient methods for the decomposition of specific circuit elements to NN connectivity have been developed [114, 115, 15].
@@ III Scalable variational circuits
The mass gap and locality of interactions appearing in physical systems with Hamiltonians such as Eq. (1) enable scaling up quantum simulations of these systems. Correlations in the ground states of such systems decay exponentially with separation beyond the correlation length (i.e., for $|i-j|>\xi,\,\langle\phi_{i}\phi_{j}\rangle\sim e^{-|i-j|/\xi}$) [118]. Here, the correlation length $\xi$ is defined to be the inverse of the gap (equivalently, the inverse of the mass of the lightest particle), $\xi=1/m_{\text{particle}}$. [^8] In the free theory, $m_{\text{particle}}\rightarrow m$ as $\delta_{\phi}\rightarrow 0$, while in the interacting theory, $m_{\text{particle}}=m_{\text{particle}}(m,\lambda)$. $m_{\text{particle}}\neq m$ in this work as a result of the digitization used $(n_{q}=2)$. This digitization may also induce interactions. Because of this, the finite-volume ground-state wavefunction for $L\gg\xi$ is exponentially close to its infinite-volume form. This, in turn, means that local observables measured in the ground state agree with their infinite-volume values up to corrections that are exponentially small in the system size, $\langle A\rangle_{L}=\langle A\rangle_{\infty}+\mathcal{O}(e^{-L/\xi})$.

### v2

@@ I Introduction
Quantum simulation holds the potential for predictive power in a wide class of fundamental physics problems that are out of reach for classical methods. In particular, quantum simulations of high-energy inelastic scattering of relativistic particles, such as those that take place in extreme cosmological events or particle colliders, promise to shed light onto the structure and properties of the Standard Model and what may lie beyond [31, 32]. Standard techniques in quantum field theory predict the outputs of scattering events in regimes where perturbation theory is valid. Although finite-volume lattice methods [33, 34] are able to compute non-perturbative quantities for specific processes (for example, see Refs. [35, 36, 37, 38, 39]), studying arbitrary scattering events is challenging. Even in the perturbative regime, conventional analytical methods only give static quantities, or results that are valid at infinitely-late times. While this approximation is valid for many quantities of interest, the dynamics of collision events are generally inaccessible. Numerical lattice techniques suffer from an exponential increase of resources due to the sign problem when attempting to simulate real-time dynamics. Furthermore, classical simulations of highly-inelastic scattering using methods based on truncations of entanglement, such as Matrix Product States (MPS), are expected to fail because of the large amount of entanglement that is believed to be generated in these events [40, 41, 42, 43, 44, 45]. As such, the dynamics of scattering of fundamental particles is a prime example of a physically-relevant problem that is challenging for classical computation, making it a promising target for quantum simulation.
@@
Much work has been done toward the goal of full-scale quantum simulations of scattering in theories and regimes where analytic results are scarce and classical computations are intractable. Proposals of wavepacket creation techniques [46, 40, 47, 48, 49, 50, 51, 52] and strategies to extract scattering observables from dynamics [53, 54] have complemented demonstrations of quantum simulations of scattering in various models [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. One of the simplest relativistic field theories in which scattering may be studied is the scalar field theory in one dimension. [^1] While it has been proven that the scalar field theory is trivial in five spacetime dimensions, and believed to be trivial in four, the existence of a non-trivial continuum limit has been rigorously established in the case of one and two spatial dimensions [67]. The continuum limit can be recovered from lattice simulations by studying systems with successively finer lattices and carefully tuning the system parameters as functions of the lattice spacing. This is the theory of coupled harmonic oscillators at each point in space with a quartic potential that governs interactions between particles. Scalar and pseudoscalar particles play key roles in many areas, including nuclear physics [68, 69], condensed matter physics [70, 71], high energy physics [72, 73, 74], as well as cosmology [75] and physics beyond the Standard Model [76, 77, 78]. Scalar fields have also been used as a venue to study phase transitions and spontaneous symmetry breaking [79, 80, 81]. Moreover, scattering in scalar field theory has been shown to be BQP-complete [82], meaning all “quantum-efficient” problems may be solved by mapping them to quantum simulations of scattering with various initial conditions. As such, this “simple” theory is a good sandbox for demonstrations of scalable quantum simulation methods. Following the seminal work of Jordan, Lee, and Preskill (JLP) [83, 84, 82], there have been numerous efforts targeting aspects of quantum simulations of scattering in scalar field theory [85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97].
@@
In this work, variational algorithms on the basis of brickwall [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108] circuits222“Brickwall”, “brickwork”, and “hardware-efficient” are commonly used to refer to dense variational ansatze using native single- and two-qubit gates. “Brickwall” is used to describe this class of circuits in this work. and other ansatze are used to simulate the scattering of two particle wavepackets on IBM’s quantum computers. While the application of variational algorithms to quantum simulations has been proposed and executed on devices in the past [109, 110, 111, 112, 113], this work serves as the first example of a quantum simulation using physics-informed brickwall circuits at scale for state preparation and time evolution. Furthermore, it is the first quantum simulation of wavepacket scattering in an interacting quantum field theory. These proof-of-concept simulations are a demonstration of the techniques that will be foundational for simulations of high-energy, high-inelasticity scattering events in more complicated theories, where rich physical phenomena are expected. The simulations proceed in several steps:
@@
By incorporating the symmetries and physical structure of the scalar field theory into the circuit design, the simulation is carried out in a scalable way, unlocking the full capabilities of state-of-the-art devices. Throughout this work, “scalable” will be used to indicate scalability for states whose structure does not change with increasing system size. Recent work has determined upper bounds of $O(10^{12})$ fault-tolerant operations required for simulations of scattering in one-dimensional scalar field theory [97]. The variational techniques to minimize circuit depth developed in this work enable approximate simulations of scattering processes with feasible gate counts for devices available today. While previous quantum simulations have taken a hybrid approach, (e.g., state preparation was done variationally, but time evolution was implemented using an algorithm with rigorous performance guarantees), this work establishes the application of physics-informed variational algorithms to all parts of the simulation. The variational methods developed in this work are expected to have broad relevance to simulating physical models of interest.
This paper is organized as follows. An overview of lattice scalar field theory in one dimension and its mapping to qubits is given in Sec. II. The algorithm for implementing scalable variational circuits (SVCs) for a class of quantum simulation problems is outlined in Sec. III. This algorithm may be seen as a general framework capable of compressing physical simulation circuits at scale. Sec. IV describes how knowledge of the physics of the theory is used to determine SVCs both for state preparation and time evolution. The errors stemming from these approximate circuit compression methods are quantified. In Sec. V, SVCs are used to simulate scattering in scalar field theory using 120 qubits of IBM’s superconducting quantum computer ibm_fez. Both the free and interacting cases are studied. A new error mitigation strategy on the basis of simulations of vacuum evolution is introduced. With this technique, the results are found to be in qualitative agreement with MPS circuit simulations, and the effect of the interaction strength on the nature of the collisions is clearly seen. Improvements and comments on the extensibility of the developed methods are presented in Sec. VI.
@@ II Lattice scalar field theory
$H$ $=\sum_{j=0}^{L-1}\frac{1}{2}\Pi_{j}^{2}+\frac{1}{2}m^{2}\phi_{j}^{2}+\frac{1}{2}(\phi_{j+1}-\phi_{j})^{2}+\frac{\lambda}{4!}\phi_{j}^{4}$ (1)
@@
Here $m$ is the bare mass, $\lambda$ is the coupling controlling the strength of $\phi^{4}$ interactions, $\Pi_{j}$ is the conjugate momentum operator obeying the canonical commutation relation $\phi_{i},\Pi_{j}]=i\delta_{ij}$, and the nearest-neighbor finite-difference operator $\nabla\phi=(\phi_{j+1}-\phi_{j})$ is used in the lattice representation of the kinetic term. The lattice spacing is set to $a=1$ in this work.
With PBCs, the eigenstates of $H$ are labeled by their spatial momenta $k=2\pi n/L$, where $n$ is an integer such that $k\in(-\pi,\pi]$. The vector of fields $\vec{\phi}=(\phi_{0},\phi_{1},\dots,\phi_{L-1})$ is related to the vector of spatial momentum eigenstates $\vec{\chi}$ by a spatial Fourier transform: $\vec{\chi}=V\vec{\phi}$, where $V=\frac{1}{\sqrt{L}}e^{i\vec{k}\cdot\vec{j}}$ is the Fourier transform matrix. In the free theory $(\lambda=0)$, the energy of a single-particle eigenstate with spatial momentum $k$ is
@@
$H=\frac{1}{L}\sum_{k}E_{k}\left(a_{k}^{\dagger}a_{k}+\frac{1}{2}\right).$ (4)
@@
Time-evolved observables in the free theory can be computed analytically by expanding the state in the appropriate basis of multiparticle excitations. For example, the time evolution of a single-wavepacket state can be written as $|\psi_{\text{wp}}\rangle(t)=\mathcal{N}\sum_{k}e^{-itE_{k}}g(k,p)a_{k}^{\dagger}|\psi_{\text{vac}}\rangle$. By rewriting $\phi^{2}_{j}$ in terms of $a_{k}$ and $a_{k}^{\dagger}$, $\langle\phi^{2}_{j}\rangle(t)$ can be obtained using Wick contractions. In the interacting $(\lambda\neq 0)$ theory, $E_{k},a_{k}^{\dagger},$ and $|\psi_{\text{vac}}\rangle$ are all functions of $\lambda$. In cases where $\frac{\lambda}{4!}$ is small, observables can be obtained by treating the $\phi^{4}$ term as a perturbation.
@@
$U_{\text{ad}}=\prod_{n=0}^{N-1}T\left\{\exp\left(-i\frac{t_{\text{ad}}}{2}\int_{\frac{(2n+1)}{2N}}^{\frac{(n+1)}{N}}H(s)ds\right)\right\}\exp\left(it_{\text{ad}}H\!\left(\frac{2n+1}{2N}\right)\right)T\left\{\exp\left(-i\frac{t_{\text{ad}}}{2}\int_{\frac{n}{N}}^{\frac{(2n+1)}{2N}}H(s)ds\right)\right\},$ (12)
where $T$ is the time-ordering operator. With a proper choice of $N$ and $t_{\text{ad}}$ depending on the parameters $m$ and $\lambda$ [83, 84, 114], it can be shown that unwanted evolution can be sufficiently suppressed.
@@ II.1 Qubit representation
At each lattice site $j$, the field takes continuous values $\phi_{j}\in(-\infty,\infty)$. Since the Hilbert space at each $j$ is infinite, the field must be digitized to be represented on a quantum computer. The JLP digitization is used to map this system onto qubits [83, 84]. Given a register of $n_{q}$ qubits, a maximum magnitude $\phi_{\text{max}}$ is chosen, [^3] There exists an optimal choice of $\phi_{\text{max}}$ for a given $n_{q}$ as a result of the Nyquist-Shannon Sampling Theorem [115, 116, 117, 118]. The optimal $\phi_{\text{max}}$ also depends on $m$ and $\lambda$, since these parameters control the support of the eigenstates of the theory (see App. A). The optimal values of $\phi_{\text{max}}$ are determined numerically for small $n_{q}$ in Ref. [85], and analytically in Refs. [119, 120]. and the field at each point in space is evenly discretized in $2^{n_{q}}$ steps of size $\delta_{\phi}$ in the range $\phi_{\text{max}},\phi_{\text{max}}]$. At each site $j$, the digitized field can take on the values
@@
$k_{\phi,j}=-\frac{\pi}{\delta_{\phi}}+\left(\ell+\frac{1}{2}\right)\frac{2\pi}{2^{n_{q}}\delta_{\phi}},\quad\ell\in[0,2^{n_{q}}-1].$ (14)
The allowed state of the field at each site is represented by qubits using a binary encoding of the $2^{n_{q}}$ allowed values from Eqs. (13) and (14):
@@
The representation of the lattice scalar field by qubits is shown in Fig. 1. This construction efficiently recovers the low-energy physics in a way that is a systematically extendable to include higher-energy contributions. In Ref. [85] it was shown that, aside from device noise and errors stemming from approximations to the time evolution operator, the error in local observables decays double-exponentially with the amount of qubits used in the site-wise digitization. [^5] That is, for a local observable, the percent error $\epsilon$ in the low-energy eigenvalues of the digitized observable compared to their undigitized counterparts scales as $\epsilon\sim 2^{-2^{n_{q}}}$. In fact, it was found in Ref. [85] that $n_{q}=4$ provides sufficient precision in local observables, so that other sources of error dominate. Furthermore, it was also found that the $\phi$ basis is advantageous over other representations, such as the basis of harmonic oscillator eigenstates, and strategies associated with improved actions that are often used in lattice QCD. This advantage stems from the simplicity and efficiency of the implementations of operators in the theory (Eqs. (16) and (17)).
The JLP prescription gives an efficient implementation of the operators using quantum gates [83, 84]. Operators in the Hamiltonian of Eq. (1) can be broken into two non-commuting groups: those that contain $\phi$, and those that contain $\Pi$. The operators made up of powers and tensor products of $\phi$ are diagonal in the $\phi$ basis, and are straightforward to implement. Aside from factors of the identity, $\phi^{2}_{j}$ and $(\phi_{j+1}-\phi_{j})^{2}$ contain only $ZZ$ operators and can be implemented using $O(n_{q}^{2})$ two-qubit gates. Similarly, $\phi^{4}_{j}$ can be implemented with $O(n_{q}^{4})$ two-qubit gates, as it has additional $ZZZZ$ terms for $n_{q}\geq 4$. $\Pi^{2}_{j}$ is translated into gates by using a site-wise quantum Fourier transform (QFT) to switch into the conjugate momentum basis where $\Pi$ is diagonal. Using this, $\Pi^{2}_{j}$ can be implemented with $O(n_{q}^{2})$ two-qubit gates as well. [^6] Note that for a device with nearest-neighbor (NN) two-qubit gate connectivity, as is the case for most superconducting quantum computers, there is an extra two-qubit gate overhead to convert circuits from arbitrary connectivity to NN. To reduce the two-qubit gate depth, efficient methods for the decomposition of specific circuit elements to NN connectivity have been developed [121, 122, 15].
@@ III Scalable variational circuits
The mass gap and locality of interactions appearing in physical systems with Hamiltonians such as Eq. (1) enable scaling up quantum simulations of these systems. Correlations in the ground states of such systems decay exponentially with separation beyond the correlation length (i.e., for $|i-j|>\xi,\,\langle\phi_{i}\phi_{j}\rangle\sim e^{-|i-j|/\xi}$) [125]. Here, the correlation length $\xi$ is defined to be the inverse of the gap (equivalently, the inverse of the mass of the lightest particle), $\xi=1/m_{\text{particle}}$. [^8] In the free theory, $m_{\text{particle}}\rightarrow m$ as $\delta_{\phi}\rightarrow 0$, while in the interacting theory, $m_{\text{particle}}=m_{\text{particle}}(m,\lambda)$. $m_{\text{particle}}\neq m$ in this work as a result of the digitization used $(n_{q}=2)$. This digitization also introduces interactions (see App. A). Because of this, the finite-volume ground-state wavefunction for $L\gg\xi$ is exponentially close to its infinite-volume form. This, in turn, means that local observables measured in the ground state agree with their infinite-volume values up to corrections that are exponentially small in the system size, $\langle A\rangle_{L}=\langle A\rangle_{\infty}+\mathcal{O}(e^{-L/\xi})$.